# Anomaly-induced vanishing of brane partition functions

### v1

@@ 1 Introduction
Consider a D-brane in weakly coupled Type II string theory, wrapping some closed submanifold $D$ of ten dimensional spacetime. A very well known result by Freed and Witten Witten:1998xy ; Freed:1999vc states that the theory of open worldsheets ending on $D$ has a global anomaly unless
@@
where $H]|_{D}\in H^{3}(D;\mathbb{Z})$ is the characteristic class of the $B$ NSNS 2-form field restricted to $D$, $N_{D}$ is the normal bundle to $D$ in the ambient space, and $W_{3}$ its third integral Stiefel-Whitney class.
The main goal of this paper is to derive a generalisation of this cancellation condition to situations where the original worldsheet derivation does not apply due to the presence of strong string coupling. Such generalisations have already been worked out in multiple cases using diverse techniques (we will provide references below), our contribution is to provide an anomaly-based argument that allows us to rederive existing generalisations in a systematic manner, and also allows us to study some cases that had not been understood before. In contrast with the original approach of Freed and Witten, our approach makes use of a (higher-form) anomaly on the worldvolume of the brane itself, and it is applicable whenever the anomaly theory for the Quantum Field Theory (QFT) on the brane is known. We emphasise that this anomaly theory is often significantly easier to understand than the worldvolume theory itself. Although in this paper we will be mostly interested in brane physics, the argument we use is in fact purely field theoretical, and we will also discuss how it applies to various QFTs of independent interest, such as the 3d minimal TQFTs Hsin:2018vcg , which we study in Section 4.1.
In order to be able to treat some of the topological subtleties that arise in the analysis we will work in the framework of differential cohomology. The main ideas of the argument can be stated without having to introduce this machinery, so we briefly summarise them here to make the logic clearer. Our approach follows from two fundamental observations. First, consider free Maxwell theory on $D$. This theory has two $\mathrm{U}(1)$ 1-form symmetries Gaiotto:2014kfa , which we denote by $\mathrm{U}(1)^{[1]}_{e}$ and $\mathrm{U}(1)^{[1]}_{m}$. The partition function of the theory coupled to a background $B_{e}$ for the electric symmetry is given (slightly schematically, we will be more precise below) by
@@
This is a close cousin of the actual theory living on a single D-brane, with the background field $B_{e}$ playing the role of the restriction of the Type II NSNS field $B$ to the brane. In this context, the Freed-Witten condition becomes the requirement that $dB_{e}]=0$ as an element of $H^{3}(D;\mathbb{Z})$. Hsieh, Tachikawa and Yonekura argued convincingly in Hsieh:2020jpj that whenever $dB_{e}]\neq 0$, the partition function of Maxwell theory should be taken to be zero. We follow this point of view, and will interpret the Freed-Witten anomaly cancellation condition as the statement that the brane partition function vanishes whenever condition (1) (or its generalisations) is not satisfied.
@@
where $\mathcal{O}(x)$ is an operator with charge $q$ under the $\mathrm{U}(1)$ symmetry, and $\Phi$ are the dynamical fields in this theory. There is a well-known argument showing that for $q\neq 0$ this expectation value vanishes on any closed [^1] In order to have a simple form of the vanishing argument, we will always assume in this paper that $\mathcal{M}^{d}$ is closed, otherwise the question of vanishing depends sensitively on the action of symmetries on the boundary conditions. manifold $\mathcal{M}^{d}$: consider a small codimension-1 sphere $S$ surrounding the insertion point $x$, and insert a symmetry generator for the $\mathrm{U}(1)$ symmetry on $S$. We denote this insertion by $U_{\alpha}(S)$, with $\alpha\in\mathbb{R}/\mathbb{Z}$ parameterising the $\mathrm{U}(1)$ transformation generated by the operator. Concretely, we have
@@
(We will motivate the terminology momentarily.) Ordinarily, introducing a symmetry generator can be seen as a modification of the background for the symmetry: since the background couples to the current via a term $A\wedge j$ in the action, if we insert a symmetry generator $U_{\alpha}(S)$ defined as above into the path integral, we are effectively shifting the background $A$ for the $\mathrm{U}(1)$ symmetry by $A\to A+2\pi\alpha\delta(S)$, with $\delta(S)$ a Dirac-delta 1-form localised on $S$. In terms of $I_{S}$, let us introduce a generalised Heaviside function $\Theta(I_{S})$ which is equal to 1 inside $I_{S}$, and 0 outside. We can then write
@@
All this, of course, is just reproducing well-known field theory phenomena in a slightly different language, but we are now in a position where we can generalise our discussion to higher symmetries, and to theories without a Lagrangian description. Let us consider first the generalisation to continuous higher-form symmetries, which is immediate. Associated to every $\mathrm{U}(1)$ $p$-form symmetry there is a conserved (dp1)(d-p-1) current $j$, satisfying $dj=q\delta^{d-p}(\Sigma^{p})$, where $\Sigma^{p}$ is the $p$-dimensional locus where we have placed charged insertions. Rigid operators are now parameterised by closed codimension $p$ submanifolds $\mathcal{N}^{d-p}$ of $\mathcal{M}^{d}$:
@@
As mentioned above, our goal in this paper is to construct a version of the vanishing argument where the net charge comes from backgrounds for the (anomalous) symmetries present in the system. We will do so in the powerful language of anomaly theories Freed:2014iua ; Monnier:2019ytc : these are invertible field theories in $d+1$ dimensions whose action reproduces the perturbative anomalies one also gets from the traditional descent procedure Zumino:1983rz , but which can be defined much more generally. In particular, anomaly theories enable us to treat non-perturbative anomalies and anomalies for discrete symmetries on an equal footing with perturbative anomalies for continuous symmetries. If we want to reproduce the anomaly descent computation in the continuous case, we can place the anomaly theory on the cylinder $\mathcal{C}\coloneqq[0,1]\times\mathcal{M}^{d}$, and then perform a gauge transformation of the background fields in the ($d+1$)-dimensional theory such that the change in the action of the anomaly theory localises on one of the endpoints of the cylinder. As a concrete example, if we are interested in a rigid transformation $e^{i\alpha}$ for some $\mathrm{U}(1)$ symmetry with background field $A$, we can perform a gauge transformation $g(t)=e^{i\alpha t}$ where $t\in[0,1]$ is the coordinate along the cylinder. This will induce a transformation $A\to A+\alpha dt$ of the background field on $\mathcal{C}$. The actions for anomaly theories are gauge-invariant up to boundary terms, so the anomalous phase will come from the boundary contributions.
@@
and $t_{\mathsf{R}}\in(0,1)$. (After regularising $\vartheta(x)$ in the standard way, this amounts to choosing a new parameterisation of $[0,1]$.) While the transformation is still pure gauge, the change in $A$ is now localised at $t=t_{\mathsf{R}}$: we have $A\to A+\alpha\delta(t-t_{\mathsf{R}})$. This changes the partition function of the anomaly theory precisely as an insertion of $\mathsf{R}_{\alpha}(\mathcal{M}^{d})$ at $t=t_{\mathsf{R}}$ would. In other words, one of the gauge transformations that we could introduce on $\mathcal{C}$ to compute the change of the phase in the partition function due to the anomaly corresponds precisely to pulling $\mathsf{R}_{\alpha}(\mathcal{M}^{d})$ into the bulk of $\mathcal{C}$.
@@
We get to choose $\lambda$ arbitrarily, as long as it is closed, so this phase factor will only be the identity for all choices of $\lambda$ if $dB_{e}]=0\in H^{3}(\mathcal{M}^{4};\mathbb{R})$. Whenever this is not true, the partition function vanishes.
@@
Finally, let us provide an alternative viewpoint on the vanishing result, which we will adopt in the rest of this paper. Take $\mathcal{C}\coloneqq[0,1]\times\mathcal{M}^{4}$ as before, and consider as an example the theory of a free Dirac fermion in four dimensions. We assume that the signature of $\mathcal{M}^{4}$ vanishes, for simplicity. As is well-known, this theory has a $\mathrm{U}(1)_{V}$ vector symmetry and a $\mathrm{U}(1)_{A}$ axial symmetry, with a mixed anomaly encoded in the presence of an $A_{A}\wedge F_{V}^{2}$ term in the anomaly theory PhysRev.177.2426 ; Bell:1969 . [^4] We are being a bit imprecise here: the anomaly theory is better described as the exponentiated $\eta$ invariant Dai:1994kq , but for the level of precision we are aiming at in this introduction the description in the text is sufficient. If we evaluate the anomaly for arbitrary backgrounds $A_{A}$, $A_{V}$ on $\mathcal{C}$ it will in general lead to a phase factor different from one. This does not imply that the partition function should vanish: if we slice $\mathcal{C}$ into constant-$t$ slices, so that we have families of backgrounds $A_{A}(t)$, $A_{V}(t)$ on $\mathcal{M}^{4}$, the partition function of the anomaly theory is telling us how the phase of the partition function changes as we move along the families of background connections. More precisely, we are studying how the phase of the determinant line bundle changes under parallel transport in the space of connections. Parallel transport on a generic bundle does indeed induce non-trivial holonomies in general. For example, if the bundle has curvature (which indicates the presence of perturbative anomalies) small loops can generate non-trivial phases. These phases do not imply that the partition function needs to vanish, they only tell us that the connection of the determinant line bundle over the space of background connections is not trivial.
@@
From this point of view, the rigid connections that we will be choosing are very special. For instance, in the ABJ case, in order to probe vanishing we could choose a connection on $\mathcal{C}$ of the form $A_{A}=\alpha dt$ and $A_{V}$ a pullback of some connection on $\mathcal{M}^{4}$ (that is, we assume $A_{V}$ to be constant on the $t$ direction, and to have no $dt$ component). If we evaluate $A_{A}\wedge F_{V}\wedge F_{V}$ on such a background it leads to an anomalous phase given by $\exp(\pi i\alpha\int_{M^{4}}F_{V}\wedge F_{V})$, which implies vanishing of the partition function whenever the index of the Dirac operator on $\mathcal{M}^{4}$ (which we recall we are assuming to have vanishing signature) is non-zero. This of course agrees with the expectation from field theory, since zero modes lead to the vanishing of the partition function, and with the classical calculation by Fujikawa 10.1093/acprof:oso/9780198529132.001.0001 . When doing the constant-$t$ slicing above, the connection on each constant-$t$ slice is always the same, and in particular $A_{A}$ vanishes. From the point of view of the determinant line bundle, we are finding that we can modify the phase of the determinant while keeping the background connection fixed.
@@
Let us finish this introduction by briefly mentioning that throughout this paper we will be dealing only with (possibly higher-form) Abelian symmetry groups. This leaves many possible directions for further work, for example the generalisation of our arguments to more general categorical symmetries, or alternatively to (possibly Abelian) subgroups of non-Abelian transformations leaving the background fixed. A recent interesting work in this last direction is Meynet:2025zem .
@@ 2 Basepoint anomaly from the mapping torus
In this section, we establish more precisely the relation between the vanishing of the partition function of a QFT and the evaluation of its anomaly on the mapping torus. As is well-known, a gauge field can carry non-trivial topological data, and an adequate language which captures all such data is differential cohomology 10.1007/BFb0075216 ; Hopkins:2002rd . The advantage of working with differential cocycles is that it makes the aforementioned relation manifest. For this purpose, we shall briefly review the notion of differential cohomology.
@@ 2.1 A brief review of differential cohomology
Suppose we have a QFT on some closed $d$-dimensional spacetime manifold $X^{d}$, and it has a $p$-form $\mathrm{U}(1)$ global symmetry Gaiotto:2014kfa . One can then couple the theory to a $(p+1)$-form background gauge field $A_{p+1}$. More precisely, let us work with differential cochains in the sense of Hopkins:2002rd . The cochain complex is defined as
@@
where $Q(\alpha,\beta)\in C^{|\alpha|+|\beta|-1}(X^{d}\times I;\mathbb{R})$ is any natural chain homotopy between the wedge product $\wedge$ for differential forms and the cup product $\cup$ for cochains 10.1007/BFb0075216 , defined to be such that
@@
where $\Lambda_{p+1}\in C^{p+1}(X^{d};\mathbb{Z})$ and $\lambda_{p}\in C^{p}(X^{d};\mathbb{R})$. There are two subclasses of this equivalence relation which are commonly known in the physics literature, namely, small gauge transformations where $\Lambda_{p+1}$ is exact and $\lambda_{p}\in C^{p}(X^{d};\mathbb{R})$ is generic, as well as large gauge transformations where $\Lambda_{p+1}\in Z^{p+1}(X^{d};\mathbb{Z})$ is closed with non-zero integral periods Diaconescu:2003bm ; Moore:2004jv ; Belov:2004ht ; Belov:2006jd ; Freed:2006yc .

### v3

@@ 1 Introduction
Consider a D-brane in weakly coupled Type II string theory, wrapping some closed submanifold $D$ of ten dimensional spacetime. A very well known result by Freed and Witten Witten (1998); Freed and Witten (1999) states that the theory of open worldsheets ending on $D$ has a global anomaly unless
@@
where $H]|_{D}\in H^{3}(D;\mathbb{Z})$ is the characteristic class of the NSNS 2-form field $B$ restricted to $D$, $N_{D}$ is the normal bundle to $D$ in the ambient space, and $W_{3}$ its third integral Stiefel-Whitney class.
The main goal of this paper is to derive a generalisation of this cancellation condition to situations where the original worldsheet derivation does not apply due to the presence of strong string coupling. Such generalisations have already been worked out in multiple cases using diverse techniques (we will provide references below), our contribution is to provide an anomaly-based argument that allows us to rederive existing generalisations in a systematic manner, and also allows us to study some cases that had not been understood before. In contrast with the original approach of Freed and Witten, our approach makes use of a (higher-form) anomaly on the worldvolume of the brane itself, and it is applicable whenever the anomaly theory for the Quantum Field Theory (QFT) on the brane is known. We emphasise that this anomaly theory is often significantly easier to understand than the worldvolume theory itself. Although in this paper we will be mostly interested in brane physics, the argument we use is in fact purely field theoretical, and we will also discuss how it applies to various QFTs of independent interest, such as the 3d minimal TQFTs Hsin et al. (2019), which we study in Section 4.1.
In order to be able to treat some of the topological subtleties that arise in the analysis we will work in the framework of differential cohomology. The main ideas of the argument can be stated without having to introduce this machinery, so we briefly summarise them here to make the logic clearer. Our approach follows from two fundamental observations. First, consider free Maxwell theory on $D$. This theory has two $\mathrm{U}(1)$ 1-form symmetries Gaiotto et al. (2015), which we denote by $\mathrm{U}(1)^{[1]}_{e}$ and $\mathrm{U}(1)^{[1]}_{m}$. The partition function of the theory coupled to a background $B_{e}$ for the electric symmetry is given (slightly schematically, we will be more precise below) by
@@
This is a close cousin of the actual theory living on a single D-brane, with the background field $B_{e}$ playing the role of the restriction of the Type II NSNS field $B$ to the brane. In this context, the Freed-Witten condition becomes the requirement that $dB_{e}]=0$ as an element of $H^{3}(D;\mathbb{Z})$. Hsieh, Tachikawa and Yonekura argued convincingly in Hsieh et al. (2022) that whenever $dB_{e}]\neq 0$, the partition function of Maxwell theory should be taken to be zero. We follow this point of view, and will interpret the Freed-Witten anomaly cancellation condition as the statement that the brane partition function vanishes whenever condition (1) (or its generalisations) is not satisfied.
@@
where $\mathcal{O}(x)$ is an operator with charge $q$ under the $\mathrm{U}(1)$ symmetry, and $\Phi$ are the dynamical fields in this theory. There is a well-known argument showing that for $q\neq 0$ this expectation value vanishes on any closed [^1] In order to have a simple form of the vanishing argument, we will always assume in this paper that $\mathcal{M}^{d}$ is closed, otherwise the question of vanishing depends sensitively on the action of symmetries on the boundary conditions. manifold $\mathcal{M}^{d}$: consider a small codimension-1 sphere $S$ surrounding the insertion point $x$, and insert a symmetry generator for the $\mathrm{U}(1)$ symmetry on $S$. We denote this insertion by $U_{\alpha}(S)$, with $\alpha\in\mathbb{R}/\mathbb{Z}$ parametrising the $\mathrm{U}(1)$ transformation generated by the operator. Concretely, we have
@@
(We will motivate the terminology momentarily.) If the symmetry is conserved, $dj$ will vanish, due to the Ward identity, away from charged operator insertions, so (6) will pick up phases precisely where any charged operators have been inserted. If the symmetry is anomalous, for example as in the case of the ABJ anomaly, $dj\propto F\wedge F$, so the background field itself makes $\mathsf{R}_{\alpha}(\mathcal{M}^{d})$ potentially non-trivial. In this paper we view the $F\wedge F$ background as providing charged insertions in the path integral.
Ordinarily, introducing a symmetry generator can be seen as a modification of the background for the symmetry: since the background couples to the current via a term $A\wedge j$ in the action, if we insert a symmetry generator $U_{\alpha}(S)$ defined as above into the path integral, we are effectively shifting the background $A$ for the $\mathrm{U}(1)$ symmetry by $A\to A+2\pi\alpha\delta(S)$, with $\delta(S)$ a Dirac-delta 1-form localised on $S$. In terms of $I_{S}$, let us introduce a generalised Heaviside function $\Theta(I_{S})$ which is equal to 1 inside $I_{S}$, and 0 outside. We can then write
@@
All this, of course, is just reproducing well-known field theory phenomena in a slightly different language, but we are now in a position where we can generalise our discussion to higher symmetries, and to theories without a Lagrangian description. Let us consider first the generalisation to continuous higher-form symmetries, which is immediate. Associated to every $\mathrm{U}(1)$ $p$-form symmetry there is a conserved (dp1)(d-p-1) current $j$, satisfying $dj=q\delta^{d-p}(\Sigma^{p})$, where $\Sigma^{p}$ is the $p$-dimensional locus where we have placed charged insertions. Rigid operators are now parametrised by closed codimension $p$ submanifolds $\mathcal{N}^{d-p}$ of $\mathcal{M}^{d}$:
@@
As mentioned above, our goal in this paper is to construct a version of the vanishing argument where the net charge comes from backgrounds for the (anomalous) symmetries present in the system. We will do so in the powerful language of anomaly theories Freed (2014); Monnier (2019): these are invertible field theories in $d+1$ dimensions whose action reproduces the perturbative anomalies one also gets from the traditional descent procedure Zumino et al. (1984), but which can be defined much more generally. In particular, anomaly theories enable us to treat non-perturbative anomalies and anomalies for discrete symmetries on an equal footing with perturbative anomalies for continuous symmetries. If we want to reproduce the anomaly descent computation in the continuous case, we can place the anomaly theory on the cylinder $\mathcal{C}\coloneqq[0,1]\times\mathcal{M}^{d}$, and then perform a gauge transformation of the background fields in the ($d+1$)-dimensional theory such that the change in the action of the anomaly theory localises on one of the endpoints of the cylinder. As a concrete example, if we are interested in a rigid transformation $e^{i\alpha}$ for some $\mathrm{U}(1)$ symmetry with background field $A$, we can perform a gauge transformation $g(t)=e^{i\alpha t}$ where $t\in[0,1]$ is the coordinate along the cylinder. This will induce a transformation $A\to A+\alpha dt$ of the background field on $\mathcal{C}$. The actions for anomaly theories are gauge-invariant up to boundary terms, so the anomalous phase will come from the boundary contributions.
@@
and $t_{\mathsf{R}}\in(0,1)$. (After regularising $\vartheta(x)$ in the standard way, this amounts to choosing a new parametrisation of $[0,1]$.) While the transformation is still pure gauge, the change in $A$ is now localised at $t=t_{\mathsf{R}}$: we have $A\to A+\alpha\delta(t-t_{\mathsf{R}})$. This changes the partition function of the anomaly theory precisely as an insertion of $\mathsf{R}_{\alpha}(\mathcal{M}^{d})$ at $t=t_{\mathsf{R}}$ would. In other words, one of the gauge transformations that we could introduce on $\mathcal{C}$ to compute the change of the phase in the partition function due to the anomaly corresponds precisely to pulling $\mathsf{R}_{\alpha}(\mathcal{M}^{d})$ into the bulk of $\mathcal{C}$.
@@
We get to choose $\lambda$ arbitrarily, as long as it is closed, so this phase factor will only be the identity for all choices of $\lambda$ if $dB_{e}]=0\in H^{3}(\mathcal{M}^{4};\mathbb{R})$. Whenever this is not true, the partition function vanishes. (The fact that the vanishing of the partition function of this theory follows from the mixed anomaly was already pointed out in Maeda and Tanizaki (2025). See also Córdova and Ohmori (2020) for applications of the same mapping torus sufficient condition for vanishing that we just reviewed to the study of gapless phases.) As we will see in the main text, a more careful analysis using differential cochains leads to the broader (in the presence of torsion) sufficient vanishing condition $dB_{e}]\neq 0\in H^{3}(\mathcal{M}^{4};\mathbb{Z})$, where $dB_{e}]$ denotes the characteristic class.
@@
Finally, let us provide an alternative viewpoint on the vanishing result, which we will adopt in the rest of this paper. Take $\mathcal{C}\coloneqq[0,1]\times\mathcal{M}^{4}$ as before, and consider as an example the theory of a free Dirac fermion in four dimensions. We assume that the signature of $\mathcal{M}^{4}$ vanishes, for simplicity. As is well-known, this theory has a $\mathrm{U}(1)_{V}$ vector symmetry and a $\mathrm{U}(1)_{A}$ axial symmetry, with a mixed anomaly encoded in the presence of an $A_{A}\wedge F_{V}^{2}$ term in the anomaly theory Adler (1969); Bell and Jackiw (1969). [^4] We are being a bit imprecise here: the anomaly theory is better described as the exponentiated $\eta$ invariant Dai and Freed (1994), but for the level of precision we are aiming at in this introduction the description in the text is sufficient. If we evaluate the anomaly for arbitrary backgrounds $A_{A}$, $A_{V}$ on $\mathcal{C}$ it will in general lead to a phase factor different from one. This does not imply that the partition function should vanish: if we slice $\mathcal{C}$ into constant-$t$ slices, so that we have families of backgrounds $A_{A}(t)$, $A_{V}(t)$ on $\mathcal{M}^{4}$, the partition function of the anomaly theory is telling us how the phase of the partition function changes as we move along the families of background connections. More precisely, we are studying how the phase of the determinant line bundle changes under parallel transport in the space of connections. Parallel transport on a generic bundle does indeed induce non-trivial holonomies in general. For example, if the bundle has curvature (which indicates the presence of perturbative anomalies) small loops can generate non-trivial phases. These phases do not imply that the partition function needs to vanish, they only tell us that the connection of the determinant line bundle over the space of background connections is not trivial.
@@
From this point of view, the rigid connections that we will be choosing are very special. For instance, in the ABJ case, in order to probe vanishing we could choose a connection on $\mathcal{C}$ of the form $A_{A}=\alpha dt$ and $A_{V}$ a pullback of some connection on $\mathcal{M}^{4}$ (that is, we assume $A_{V}$ to be constant on the $t$ direction, and to have no $dt$ component). If we evaluate $A_{A}\wedge F_{V}\wedge F_{V}$ on such a background it leads to an anomalous phase given by $\exp(\pi i\alpha\int_{M^{4}}F_{V}\wedge F_{V})$, which implies vanishing of the partition function whenever the index of the Dirac operator on $\mathcal{M}^{4}$ (which we recall we are assuming to have vanishing signature) is non-zero. This of course agrees with the expectation from field theory, since zero modes lead to the vanishing of the partition function, and with the classical calculation by Fujikawa Fujikawa and Suzuki (2004). When doing the constant-$t$ slicing above, the connection on each constant-$t$ slice is always the same, and in particular $A_{A}$ vanishes. From the point of view of the determinant line bundle, we are finding that we can modify the phase of the determinant while keeping the background connection fixed.
@@
Let us finish this introduction by briefly mentioning that throughout this paper we will be dealing only with (possibly higher-form) Abelian symmetry groups. This leaves many possible directions for further work, for example the generalisation of our arguments to more general categorical symmetries, or alternatively to (possibly Abelian) subgroups of non-Abelian transformations leaving the background fixed. A recent interesting work in this last direction is Meynet et al. (2025).
@@ 2 Basepoint anomaly from the mapping torus
In this section, we establish more precisely the relation between the vanishing of the partition function of a QFT and the evaluation of its anomaly on the mapping torus. As is well-known, a gauge field can carry non-trivial topological data, and an adequate language which captures all such data is differential cohomology Cheeger and Simons (1985); Hopkins and Singer (2005). The advantage of working with differential cocycles is that it makes the aforementioned relation manifest. For this purpose, we shall briefly review the notion of differential cohomology.
@@ 2.1 A brief review of differential cohomology
Suppose we have a QFT on some closed $d$-dimensional spacetime manifold $X^{d}$, and it has a $p$-form $\mathrm{U}(1)$ global symmetry Gaiotto et al. (2015). One can then couple the theory to a $(p+1)$-form background gauge field $A_{p+1}$. More precisely, let us work with differential cochains in the sense of Hopkins and Singer (2005). The cochain complex is defined as
@@
where $Q(\alpha,\beta)\in C^{|\alpha|+|\beta|-1}(X^{d}\times I;\mathbb{R})$ is any natural chain homotopy between the wedge product $\wedge$ for differential forms and the cup product $\cup$ for cochains Cheeger and Simons (1985), defined to be such that
@@
where $\Lambda_{p+1}\in C^{p+1}(X^{d};\mathbb{Z})$ and $\lambda_{p}\in C^{p}(X^{d};\mathbb{R})$. There are two subclasses of this equivalence relation which are commonly known in the physics literature, namely, small gauge transformations where $\Lambda_{p+1}$ is exact and $\lambda_{p}\in C^{p}(X^{d};\mathbb{R})$ is generic, as well as large gauge transformations where $\Lambda_{p+1}\in Z^{p+1}(X^{d};\mathbb{Z})$ is closed with non-zero integral periods Diaconescu et al. (2003b); Moore (2005); Belov and Moore (2004, 2006); Freed et al. (2007).